Give the steps you will use to separate the variable and then solve the equation:
3n- 2 = 46

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

First we have to add 2 to the both sides of the equation,
Then, we get,
= 3n–2 + 2 = 46 + 2
= 3n = 48
Now,
We have to divide both sides of the equation by 3,
Then, we get,
= 3n/3 = 48/3
= n = 16

Thanks